# MCI-GRU: Stock Prediction Model Based on Multi-Head Cross-Attention and Improved GRU

### v1

@@ 1 Introduction
Time series models [5], such as GRU and LSTM [6], have been widely utilized in stock prediction due to their significant advantages in capturing temporal dependencies within sequential data. However, these models exhibit limitations when addressing long-term dependencies in financial markets. Long-term trends and large-scale fluctuations are often obscured by noise, making it challenging for these models to extract valuable long-term dependency information from such noisy data effectively. Additionally, financial markets are characterized by a high degree of nonlinearity, with rapid shifts in market behavior driven by changes in investor sentiment, unexpected events, and other factors. These models often lack sufficient sensitivity and flexibility in handling such nonlinearity and abrupt events. Furthermore, they face challenges in the flexible selection and effective utilization of critical historical information. Given the vast and disorganized nature of data in financial markets, identifying the most relevant features for prediction has become a critical issue.
@@
Artificial intelligence technologies’ rapid advancement [9], particularly in Graph Neural Networks (GNNs) [10] and Reinforcement Learning (RL) [11], has introduced unprecedented potential for stock prediction. These technologies, through innovative algorithmic designs and deep learning models, promise to enhance the ability to capture the complex dynamics of financial markets. For example, methods that employ GNNs are able to precisely capture the complex and diverse interdependencies within financial data by modeling the relationships between stocks as graph structures. This approach not only uncovers deep connections that are difficult for traditional models to reveal but also more effectively reflects the nonlinear characteristics of the market. Meanwhile, methods based on RL progressively learn and optimize trading strategies by simulating continuous interaction with the market environment. This method is highly adaptive, enabling dynamic strategy adjustments to cope with the market’s rapidly changing conditions. However, despite the considerable promise of these emerging methodologies, they still face significant challenges in practical application. Firstly, these models often rely on large-scale, high-quality datasets, which are difficult to acquire or construct in real-world scenarios. When confronted with data sparsity and noise—common issues in financial markets—the predictive performance of these models can be severely compromised. Additionally, the computational complexity of these methods is high; the training process is not only time-consuming and resource-intensive but also demands substantial computational power, significantly limiting their widespread adoption in practical financial applications. More critically, a fundamental limitation of current models lies in their inability to capture unobservable latent market states effectively. Factors such as market sentiment, investor expectations, microstructural elements, and participant behavior patterns play crucial roles in shaping market dynamics. However, the failure to adequately account for these latent factors often leads to a superficial understanding of the market, thereby constraining the predictive accuracy and practical utility of these models.
@@ 2.2 Deep Learning and Reinforcement Learning Methods
Deep learning models often demonstrate instability when confronted with extreme market fluctuations, such as those experienced during the 2008 financial crisis [39] and the 2019 COVID-19 pandemic [40]. In response, reinforcement learning models have gained prominence due to their adaptability and capacity for continuous learning. Reinforcement learning approaches in investment strategies can be broadly categorized into two types: value-based and policy-based [41, 42]. Value-based approaches involve learning a critique to estimate the expected outcomes of trading actions within the market. Common value-based approaches in investment strategies include Q-learning [43], Deep Q-learning [44, 45], Recurrent Reinforcement Learning [46, 47], and Sarsa [48, 49]. However, a major limitation of value-based approaches is the complexity of accurately approximating the market environment with a critic. As a result, policy-based approaches [50, 51] are often regarded as more suitable for financial markets. For example, [52] integrated deep attention networks with reinforcement learning, optimizing parameters through discrete agent actions to maximize the Sharpe ratio of investments. To further balance profit and risk, [53] introduced a multi-objective deep reinforcement learning (MODRL) approach for intraday trading of stock index futures, combining deep learning for feature extraction with reinforcement learning for decision-making. Despite their potential, reinforcement learning models face challenges, such as the requirement for large datasets and difficulties in model interpretability, which can hinder their practical application in financial markets.
@@ 2.3 Graph Neural Networks and Latest Methods
In recent years, Graph Neural Networks [54, 55, 56] have garnered significant attention in stock prediction due to their capacity to capture complex interdependencies within financial data. For instance, [57] proposed a hybrid model that integrates Recurrent Neural Networks with GNN, facilitating real-time predictions. Similarly, [58] introduced a hierarchical attention mechanism into GNN, thereby improving the model’s ability to analyze multi-level market dependencies and perform structured analyses of stock trends. To capture a stock’s intrinsic value more accurately, [59] developed a Higher-order Graph Attention Network (H-GAT), which differentiates itself by modeling complex subgraph structures involving more than two stocks and incorporating both technical and fundamental factors. This approach contrasts with traditional GNNs, which typically consider only simple pairwise relationships, thus enhancing the model’s ability to reflect the intrinsic value of stocks. However, many graph-based models frequently overlook the diversity of stock price changes and the temporal dynamics inherent in these fluctuations, necessitating the development of more innovative graph-based approaches. For instance, [60] introduced a market-guided stock transformer that can dynamically simulate the instantaneous and cross-temporal correlations among stocks, leading to enhanced accuracy in stock trend predictions. Additionally, [61] successfully integrated long-term trends, short-term fluctuations, and sudden events into a cohesive graph-based framework, significantly surpassing traditional methods by accounting for the multi-scale nature of market dynamics. Despite their advantages in stock prediction, Graph Neural Networks exhibit several limitations. These include insufficient modeling capabilities for complex nonlinear relationships and anomalous scenarios within the stock market, as well as a lack of robustness to data sparsity and noise. Furthermore, certain graph-based models may be prone to overfitting, particularly in contexts characterized by limited or imbalanced training data.
@@ 3.2.2 Introduction of the Improved GRU
where $\alpha_{t}\in\mathbb{R}$ is the attention weight vector. The core idea of the attention mechanism is to allocate attention weights by calculating the similarity between the query, key, and value. In this model, the hidden state $h_{t-1}\in\mathbb{R}^{d_{h}}$ from the previous time step is treated as the query, while the current time step’s input $x_{t}$ serves as both the key and value. The attention calculation involves the following steps:
@@ 3.2.3 Update of the Hidden State
With the new reset gate $r_{t}^{{}^{\prime}}$, the update formula for the GRU’s hidden state is adjusted as follows:
@@ 3.4.1 Initialization of Market Latent States
    Number of Latent States ($d_{r}$): The model initializes $d_{r}$ latent state vectors. The number $d_{r}$ is a hyperparameter that can be tuned based on the complexity of the market being modeled. A larger $d_{r}$ allows the model to capture a greater variety of latent factors.
@@
    Dimensionality ($d_{i}$): Each latent state vector is of dimensionality $d_{i}$, which should match the dimensionality of the stock features that the model has learned so far. This ensures compatibility when these latent states interact with the outputs from previous layers.
The initialized latent state vectors are denoted as $R_{1}$ and $R_{2}$, corresponding to the two main types of information the model processes: time series information (from the Improved GRU output $A_{1}$) and cross-sectional information (from the GAT output $A_{2}$). The dimensions of $R_{1}$ and $R_{2}$ are both $(d_{r},d_{i})$.
@@ 3.4.2 Multi-head Cross-Attention Mechanism
Multi-head Cross-Attention between $R_{1}$ and $A_{1}$: In the multi-head cross-attention mechanism, the interaction between $A_{1}$ and $R_{1}$ can be modeled through the cross-attention process, where $A_{1}$ is used as the query and $R_{1}$ is used as the key and value. Here’s how the multi-head cross-attention mechanism operates between $A_{1}$ and $R_{1}$:
@@ 3.5 Model Prediction and Loss Calculation Layer
The Loss Calculation Layer is the final stage of the model, responsible for synthesizing the outputs from previous layers and generating predictions. This layer also defines how the model is trained by computing the difference between the predicted values and the ground truth, which is then minimized during the training process.
@@ 4.1.1 Datasets
We structure our datasets following a temporal sequence, dividing them into distinct phases for training (from January 1, 2018, to December 31, 2021), validation (from January 1, 2022, to December 31, 2022), and testing (from January 1, 2023, to December 31, 2023). In our forecasting approach, we employe features derived from the previous 60 trading days to predict stock return rankings over the next 21 trading days. This methodology closely mirrors the decision-making process in real-world trading scenarios. For the baseline analysis, we utilize data from the four aforementioned stock markets, concentrating on six key financial indicators: open price, close price, high price, low price, turnover, and volume. We commence by implementing procedures for outlier detection and normalization to ensure data integrity and reduce the impact of anomalous values. Following this, we calculate the daily return for each stock as the label during training, defined as the percentage change between the closing prices of consecutive trading days.
@@ 4.1.3 Parameter Settings
We configure the time window $t$ to encompass the preceding 10 days of historical stock data as the training input for the model. The model architecture comprises four primary components, each designed to capture distinct patterns within the financial time series.
In the “Use Improved GRU to Capture Temporal Features” module, we incorporate an attention mechanism to replace the conventional reset gate in the GRU framework. This module consists of two layers, with the first layer containing 32 neurons and the second layer comprising 10 neurons. The module’s output, denoted as $A_{1}$, is responsible for extracting temporal dependencies from the stock data across the historical window.
The “Use GAT to Capture Cross-sectional Features” is structured with two layers, each with tunable parameters. The initial GAT layer consists of 32 neurons and employs 4 attention heads, while the second GAT layer reduces the feature dimensionality to 4 neurons. The relationships between nodes in the graph, representing stocks, are defined based on the correlation of stock returns over the previous year. A threshold of 0.8 is applied to filter stock pairs with significant correlations during training. The output from the GAT layer, denoted as $A_{2}$, captures the structural dependencies inherent in the stock correlation graph.
In the “Use Multi-head Cross-attention to Capture Latent State Feature”, two learnable market hidden state vectors, $R_{1}$ and $R_{2}$, each of dimension 32, are initialized. A multi-head cross-attention mechanism is employed, where $R_{1}$ interacts with $A_{1}$ and $R_{2}$ interacts with $A_{2}$, resulting in representation vectors $B_{1}$ and $B_{2}$, respectively. The head of the multi-head cross-attention is $4$. These representation vectors encapsulate the latent market states by leveraging both temporal and structural characteristics derived from the stock data.
The “Model Prediction and Loss Calculation Layer” integrates the outputs $A_{1}$, $A_{2}$, $B_{1}$ and $B_{2}$, which are concatenated and passed through a two-layer GAT for final prediction. This final module is responsible for forecasting future stock trends by fusing both temporal and relational information.
The model is trained with a batch size of 32. The mean squared error (MSE) is adopted as the loss function, and the Adam optimizer is utilized with an initial learning rate of 0.0002. For practical application, each trading day a virtual investment portfolio is constructed, comprising the top 10 stocks ranked by the model’s predicted returns.

### v2

@@ 1 Introduction
Time series models [5, 6], such as GRU and LSTM [7], have been widely utilized in stock prediction due to their significant advantages in capturing temporal dependencies within sequential data. However, these models exhibit limitations when addressing long-term dependencies in financial markets. Long-term trends and large-scale fluctuations are often obscured by noise, making it challenging for these models to extract valuable long-term dependency information from such noisy data effectively. Additionally, financial markets are characterized by a high degree of nonlinearity, with rapid shifts in market behavior driven by changes in investor sentiment, unexpected events, and other factors. These models often lack sufficient sensitivity and flexibility in handling such nonlinearity and abrupt events. Furthermore, they face challenges in the flexible selection and effective utilization of critical historical information. Given the vast and disorganized nature of data in financial markets, identifying the most relevant features for prediction has become a critical issue.
@@
Artificial intelligence technologies’ rapid advancement [10], particularly in Graph Neural Networks (GNNs) [11, 12] and Reinforcement Learning (RL) [13], has introduced unprecedented potential for stock prediction. These technologies, through innovative algorithmic designs and deep learning models, promise to enhance the ability to capture the complex dynamics of financial markets. For example, methods that employ GNNs are able to precisely capture the complex and diverse interdependencies within financial data by modeling the relationships between stocks as graph structures. This approach not only uncovers deep connections that are difficult for traditional models to reveal but also more effectively reflects the nonlinear characteristics of the market. Meanwhile, methods based on RL progressively learn and optimize trading strategies by simulating continuous interaction with the market environment. This method is highly adaptive, enabling dynamic strategy adjustments to cope with the market’s rapidly changing conditions. However, despite the considerable promise of these emerging methodologies, they still face significant challenges in practical application. Firstly, these models often rely on large-scale, high-quality datasets, which are difficult to acquire or construct in real-world scenarios. When confronted with data sparsity and noise—common issues in financial markets—the predictive performance of these models can be severely compromised. Additionally, the computational complexity of these methods is high; the training process is not only time-consuming and resource-intensive but also demands substantial computational power, significantly limiting their widespread adoption in practical financial applications. More critically, a fundamental limitation of current models lies in their inability to capture unobservable latent market states effectively. Factors such as market sentiment, investor expectations, microstructural elements, and participant behavior patterns play crucial roles in shaping market dynamics. However, the failure to adequately account for these latent factors often leads to a superficial understanding of the market, thereby constraining the predictive accuracy and practical utility of these models.
@@ 2.2 Deep Learning and Reinforcement Learning Methods
Deep learning models often demonstrate instability when confronted with extreme market fluctuations, such as those experienced during the 2008 financial crisis [41] and the 2019 COVID-19 pandemic [42]. In response, reinforcement learning models [43] have gained prominence due to their adaptability and capacity for continuous learning. Reinforcement learning approaches in investment strategies can be broadly categorized into two types: value-based and policy-based [44, 45]. Value-based approaches involve learning a critique to estimate the expected outcomes of trading actions within the market. Common value-based approaches in investment strategies include Q-learning [46], Deep Q-learning [47, 48], Recurrent Reinforcement Learning [49, 50], and Sarsa [51, 52]. However, a major limitation of value-based approaches is the complexity of accurately approximating the market environment with a critic. As a result, policy-based approaches [53, 54] are often regarded as more suitable for financial markets. For example, [55] integrated deep attention networks with reinforcement learning, optimizing parameters through discrete agent actions to maximize the Sharpe ratio of investments. To further balance profit and risk, [56] introduced a multi-objective deep reinforcement learning (MODRL) approach for intraday trading of stock index futures, combining deep learning for feature extraction with reinforcement learning for decision-making. Despite their potential, reinforcement learning models face challenges, such as the requirement for large datasets and difficulties in model interpretability, which can hinder their practical application in financial markets.
@@ 2.3 Graph Neural Networks and Latest Methods
In recent years, Graph Neural Networks [57, 58, 59] have garnered significant attention in stock prediction due to their capacity to capture complex interdependencies within financial data. For instance, [60] proposed a hybrid model that integrates Recurrent Neural Networks with GNN, facilitating real-time predictions. Similarly, [61] introduced a hierarchical attention mechanism into GNN, thereby improving the model’s ability to analyze multi-level market dependencies and perform structured analyses of stock trends. To capture a stock’s intrinsic value more accurately, [62] developed a Higher-order Graph Attention Network (H-GAT), which differentiates itself by modeling complex subgraph structures involving more than two stocks and incorporating both technical and fundamental factors. This approach contrasts with traditional GNNs, which typically consider only simple pairwise relationships, thus enhancing the model’s ability to reflect the intrinsic value of stocks. However, many graph-based models frequently overlook the diversity of stock price changes and the temporal dynamics inherent in these fluctuations, necessitating the development of more innovative graph-based approaches. For instance, [63] introduced a market-guided stock transformer that can dynamically simulate the instantaneous and cross-temporal correlations among stocks, leading to enhanced accuracy in stock trend predictions. Additionally, [64] successfully integrated long-term trends, short-term fluctuations, and sudden events into a cohesive graph-based framework, significantly surpassing traditional methods by accounting for the multi-scale nature of market dynamics. [65] proposed ECHO-GL, a model that leverages earnings call-derived relations and multimodal graph learning to predict stock movements and generate temporal price trajectories. [66] proposed the Multi-relational Dynamic Graph Neural Network (MDGNN), which integrates discrete dynamic graph modeling with Transformer-based temporal encoding to capture evolving stock relationships, demonstrating superior performance over state-of-the-art methods in stock movement prediction. [67] proposed the DANSMP model for stock movement prediction, which utilizes a Market Knowledge Graph comprising bi-typed entities and hybrid relations, combined with Dual Attention Networks to effectively capture momentum spillover signals and enhance predictive accuracy. Despite their promising capabilities in stock prediction, GNNs exhibit several notable limitations. These models often struggle to effectively capture complex nonlinear relationships and account for anomalous market scenarios, which are crucial in financial forecasting. Moreover, GNN-based models tend to be sensitive to data sparsity and noise, which can undermine their robustness, especially in real-world financial datasets where missing or unreliable data is common. Additionally, certain graph-based approaches, such as those proposed by [65], [66], and [67], may face challenges in generalization due to their reliance on specific data structures and sophisticated attention mechanisms, which can lead to overfitting, particularly in environments with limited or imbalanced training data. These models, while effective in their respective domains, often struggle with scalability and flexibility when applied to diverse market conditions and broader datasets, limiting their applicability in more volatile or unpredictable markets. Thus, enhancing the robustness and generalizability of graph-based models remains a key area for improvement in stock movement prediction.
@@ 3.2.2 Introduction of the Improved GRU
where $\alpha_{t}\in\mathbb{R}$ is the attention weight vector. The core idea of the attention mechanism is to allocate attention weights by calculating the similarity between the query, key, and value. In this model, the hidden state $h_{t-1}\in\mathbb{R}^{d_{h}}$ from the previous time step is treated as the query, while the current time step’s input $x_{t}$ serves as both the key and value. Attention calculation includes three parts: Linear Transformations of Query, Key, and Value, Weighted Sum, and Attention Weight Calculation. The specific implementation and mathematical formula are as follows:
@@ 3.2.3 Update of the Hidden State
The calculation of the update gate $z_{t}$ remains unchanged. With the new reset gate $r_{t}^{{}^{\prime}}$, the update formula for the GRU’s hidden state is adjusted as follows:
@@ 3.4.1 Initialization of Market Latent States
    Number of Latent States ($d_{r}$): The model initializes $d_{r}$ latent state vectors, where $d_{r}$ is a hyperparameter. This parameter determines the number of latent states, and it is typically adjusted based on the complexity of the modeled market. A larger $d_{r}$ allows the model to capture a wider variety of potential market factors, thereby enhancing its representational capacity. However, a $d_{r}$ that is too large could increase computational costs and introduce the risk of overfitting. Therefore, selecting an appropriate $d_{r}$ value is crucial, and it is usually obtained through cross-validation or experimental adjustments to achieve the best performance. In this paper, we set it to $16$.
@@
    Latent State Dimension ($d_{i}$): The dimension of each latent state vector is $d_{i}$, which should match the dimensionality of the stock features learned by the model in previous layers such as the GRU and GAT layers. Ensuring that $d_{i}$ is consistent with the output dimensions of the earlier layers allows these latent state vectors to interact effectively with other layers. The choice of $d_{i}$ typically depends on the needs of the model and the characteristics of the data. A larger $d_{i}$ provides more representational power but may increase computational complexity. We set it to $32$ In this paper.
The initialized latent state vectors are denoted as $R_{1}\in\mathbb{R}^{d_{h}\times d_{i}}$ and $R_{2}\in\mathbb{R}^{d_{h}\times d_{i}}$, corresponding to the two main types of information processed by the model: time series information (from the improved GRU output $A_{1}$ and cross-sectional information (from the GAT output $A_{2}$. These two types of information play different roles in the model, ensuring that the model can fully capture the dynamic changes in the market. The dimensions of $R_{1}$ and $R_{2}$ are both $(d_{r},d_{i})$, ensuring their representational capacity and enabling effective interaction with subsequent network layers.
@@ 3.4.2 Multi-head Cross-Attention Mechanism
Multi-head Cross-Attention between $R_{1}$ and $A_{1}$: In the multi-head cross-attention mechanism, the interaction between $A_{1}$ and $R_{1}$ can be modeled through the cross-attention process, where $A_{1}$ is used as the query and $R_{1}$ is used as the key and value. Using $A_{1}$ as the query and $R_{1}$ as both the key and value follows the natural logic of information processing, where $A_{1}$ guides the update of $R_{1}$, allowing effective integration of time series information into the latent state. Using $R_{1}$ as the query is logically flawed for two reasons: the information flow is reversed, as $A_{1}$ should influence $R_{1}$, and $R_{1}$, being randomly initialized, lacks a clear objective as a query target. The goal is to refine $R_{1}$ using $A_{1}$, not the other way around. Here’s how the multi-head cross-attention mechanism operates between $A_{1}$ and $R_{1}$:
@@
  • 4.
    Calculation and update process of $R_{1}$: During the model training process, the update of $R_{1}$ is achieved by calculating the gradient of the loss function with respect to $R_{1}$ through backpropagation. First, the model generates predictions through a forward pass and calculates the loss between the predictions and the actual values. Common loss functions include Mean Squared Error. Through the backpropagation algorithm, we can compute the gradient of the loss function $L$ with respect to $R_{1}$, i.e., $\frac{\partial L}{\partial R_{1}}$. Since $R_{1}$ affects the output of subsequent layers, the gradient of the loss function is propagated back to $R_{1}$ through the chain rule. During backpropagation, the gradients of the loss function with respect to $K_{i}$ and $V_{i}$ will propagate back to $R_{1}$, and the gradient computation is as follows:
    $\centering\begin{split}\frac{\partial L}{\partial R_{1}}=\sum_{i=1}^{h^{\prime}}\frac{\partial L}{\partial K_{i}}\cdot\frac{\partial K_{i}}{\partial R_{1}}\\ \frac{\partial L}{\partial R_{1}}=\sum_{i=1}^{h^{\prime}}\frac{\partial L}{\partial V_{i}}\cdot\frac{\partial V_{i}}{\partial R_{1}}\end{split}\centering\@add@centering\@add@centering$ (18)
    In practice, the gradient is propagated through the keys ($K_{i}$) and values ($V_{i}$) of each attention head and finally aggregated back to $R_{1}$. The gradient is computed based on the linear dependencies $K_{i}=R_{1}W_{i}^{K}$ and $V_{i}=R_{1}W_{i}^{V}$. Once the gradient of the loss function with respect to $R_{1}$ is calculated, the model updates $R_{1}$ using the gradient descent algorithm. The update rule is:
    $\centering\begin{split}R_{1}=R_{1}-\eta\cdot\frac{\partial L}{\partial R_{1}}\end{split}\centering\@add@centering\@add@centering$ (19)
    where $\eta$ is the learning rate, controlling the step size of the parameter update. Through this process, $R_{1}$ is adjusted according to the gradient calculated by backpropagation, allowing it to more effectively represent the latent market states and capture the complex patterns in the data.
@@ 3.5 Model Prediction and Loss Calculation Layer
The Loss Calculation Layer is the final stage of the model and is responsible for synthesizing the outputs from previous layers and generating predictions. This layer also defines how the model is trained by computing the difference between the predicted values and the ground truth, which is then minimized during the training process.
@@ 3.5.2 Final Prediction with GAT Layers
Although both GATs are used, their roles differ. The first GAT focuses on extracting cross-sectional information, capturing static dependencies between stocks based on historical correlations, and generating $A_{2}$. The second GAT, with input $Z$ (a fusion of multiple features), handles multimodal information fusion, captures higher-order relationships, and refines features. It combines information from different sources, learns complex correlations like indirect and dynamic relationships, and performs dimensionality reduction for more compact and efficient feature representations for final prediction.
@@ 4.1.1 Datasets
We structure our datasets following a temporal sequence, dividing them into distinct phases for training (from January 1, 2018, to December 31, 2021), validation (from January 1, 2022, to December 31, 2022), and testing (from January 1, 2023, to December 31, 2023). In our forecasting approach, we employ features derived from the previous 60 trading days to predict stock return rankings over the next 21 trading days. This methodology closely mirrors the decision-making process in real-world trading scenarios. For the baseline analysis, we utilize data from the four aforementioned stock markets, concentrating on six key financial indicators: open price, close price, high price, low price, turnover, and volume. We commence by implementing procedures for outlier detection and normalization to ensure data integrity and reduce the impact of anomalous values. Following this, we calculate the daily return for each stock as the label during training, defined as the percentage change between the closing prices of consecutive trading days.
@@ 4.1.2 Evaluation Metrics
Table 2: Experimental parameter summary table.
Module Name Parameter Name Value
Use Improved GRU to Capture Temporal Features number of GRU layers 2
number of neurons in the first layer 32
number of neurons in the second layer 10
Use GAT to Capture Cross-sectional Features number of GAT layers 2
number of neurons in the first layer 32
number of neurons in the second layer 4
number of attention heads 4
threshold for connecting stock nodes $judge_{value}$ 0.8
Use Multi-head Cross-attention to Capture Latent State Feature number of latent states $d_{r}$ 32
dimension of each latent state $d_{i}$ 16
Model Prediction and Loss Calculation Layer number of GAT layers 1
number of neurons in the first layer 32
Training Process batch size 32
loss function MSE
optimizer adam
learning rate 0.0002
@@ 4.1.3 Parameter Settings
We configure the time window $t$ to include the historical stock data from the previous 10 days as the model’s input for training. The model architecture consists of four primary components, each designed to capture different patterns in financial time series. Each of these components contains its own set of parameters, as summarized in Table 2.
In the “Use Improved GRU to Capture Temporal Features” module, we incorporate an attention mechanism to replace the traditional reset gate in the GRU framework. This module consists of two layers: the first layer contains 32 neurons, and the second layer contains 10 neurons. The output of this module is represented as $A_{1}$, which is responsible for extracting temporal dependencies from the stock data within the historical window. In the “Use GAT to Capture Cross-sectional Features” module, the architecture consists of two layers, each with tunable parameters. The number of attention heads is a key factor affecting the model’s expressive capability, and it was fine-tuned through experiments. We tested four values: 2, 4, 6, and 8, and ultimately chose 4. The initial GAT layer consists of 32 neurons and 4 attention heads, while the second GAT layer reduces the feature dimension to 4 neurons. The relationships between the stock nodes in the graph are defined based on the correlation of stock returns from the previous year. During training, a threshold value of $judge_{value}=0.8$ is used to filter stock pairs with significant correlations. The $judge_{value}$ parameter was optimized through experimental tuning, testing values of 0.2, 0.4, 0.6, 0.8, and 1.0, with 0.8 being the final choice. The output of the GAT layer is represented as $A_{2}$, which captures the structural dependencies inherent in the stock correlation graph. In the “Use Multi-head Cross-attention to Capture Latent State Feature” module, two learnable market latent state vectors, $R_{1}\in\mathbb{R}^{d_{h}\times d_{i}}$ and $R_{2}\in\mathbb{R}^{d_{h}\times d_{i}}$, are initialized, both with dimensions $(32,16)$. The choice of the hyperparameters $d_{r}$ (number of latent states) and $d_{i}$(dimension of each latent state) is typically influenced by the complexity of the data and the task requirements. We optimized these parameters through grid search to balance computational efficiency and model expressiveness. We tested $d_{r}$ values of 8, 16, 24, 32, and 40, ultimately selecting 32. For $d_{i}$, we tested values of 4, 8, 12, 16, and 20, with 16 being chosen as the final value. The multi-head cross-attention mechanism involves interactions between $R_{1}$ and $A_{1}$, as well as $R_{2}$ and $A_{2}$, producing representation vectors $B_{1}$ and $B_{2}$. With 4 attention heads, the multi-head cross-attention mechanism encapsulates latent market states by leveraging temporal and structural features derived from the stock data. Finally, the “‘Model Prediction and Loss Calculation Layer” integrates the outputs $A_{1}$, $A_{2}$, $B_{1}$, and $B_{2}$. These outputs are concatenated and passed through one additional layers of GAT for final prediction. This final module is responsible for predicting future stock trends by merging both temporal and relational information.
The batch size of 32 is chosen based on experimental tuning. While 32 is a commonly used value and works well for most models, we acknowledge that it can influence the stability and efficiency of the training process. We selected 32 as a compromise, ensuring efficient utilization of computational resources while avoiding excessively large variance during training. Future research may further explore different batch sizes through experiments to optimize performance for specific tasks and datasets. For the loss function, we used Mean Squared Error (MSE), which is the most common choice for regression tasks, particularly when the target variable is continuous. We opted for MSE because it provides stable performance in most regression tasks and effectively measures the squared error of predictions. The Adam optimizer was chosen due to its strong performance across various deep learning tasks, particularly its adaptability in the optimization process, which allows for good convergence with minimal tuning. The initial learning rate of 0.0002 was determined through preliminary experiments, with smaller learning rates typically ensuring stable convergence. This value was selected based on its performance in our experiments but can be adjusted depending on the specific task at hand. For practical application, each trading day a virtual investment portfolio is constructed, comprising the top 10 stocks ranked by the model’s predicted returns.